The order reads: 900mL NS IV 5 hr. Calculate the f
Answer #1

1. Volume = 900 ml

Time = 5hr = 300 min

Drop factor = 10 drops per ml

Flow rate = 10 * 900/300 = 30 drops per min = 90 ml per hour [ 20 drops = 1 ml]

2. convert 80 microdrops per minute to ml per hr

80 microdrops per min = 0.00008*60/20 = 0.00024 ml per hour

3. volume = 1000 ml

time = 8 hrs = 480 min

drop factor = 20 gtt/ml

Flow rate = 1000 * 20 / 480 = 41.667 gtt/min

4. volume = 300 ml

flow rate = 60 ml/h

time for infusion of 5% D/W = 300/60 = 5 hours